OpenStudy (anonymous):

Nick is solving the equation 3x2 = 20 – 7x with the quadratic formula. What values should he use for a, b, and c? (Points : 1) a = 3 b = 20 c = –7 a = 3 b = –7 c = 20 a = 3 b = 7 c = –20 a = 3 b = –20 c = 7

OpenStudy (welshfella):

write the eqaution in the form ax^2 + bx + c = 0 3x^2 = 20 - 7x 3x^2 + 7x - 20 = 0

OpenStudy (welshfella):

now compare the first equation term by term with the last eg a = 3 what are b and c?

OpenStudy (welshfella):

do you understand what i mean?

OpenStudy (welshfella):

are you there?

OpenStudy (anonymous):

sorry i was doing something im back

OpenStudy (welshfella):

ok do you follow the above ?

OpenStudy (anonymous):

no not really honestly i have about 10 to 15 more test to take in algebra today.... i just need the answers on the ones i don't understand....

OpenStudy (welshfella):

I cant just give answers i'm afraid i'll just draw this to see if you might get it